OpenStudy (anonymous):

When an altitude is drawn to the hypotenuse of a right triangle, the lengths of the segments of the hypotenuse are 16 and 25. What is the length of the altitude? 41 50 25 20

OpenStudy (anonymous):

help me please @heybhai

OpenStudy (mathstudent55):

It's the geometric mean of the two lengths.

OpenStudy (anonymous):

 this is a shortcut method remember whenever u writing a competitive exam . DB^2=CD*DA altitude=sqrt(16*25) =4*5 =20